# Quantifying Phase Unbalance and Coordination Impacts on Distribution Network Flexibility

### v1

@@ Abstract
The increasing integration of distributed energy resources (DER) provides distribution system operators (DSO) with new flexible resources to support more efficient operation and planning of distribution networks. To utilise these resources, various DER flexibility aggregation methods have been proposed in the literature, such as aggregated P-Q flexibility areas at the interface with other networks. However, whereas focusing on estimating the limits of flexibility services, existing studies make the critical assumption that all available flexible units are perfectly coordinated to jointly provide flexibility and manage network constraints. Moreover, due to the extensive use of single-phase power flow analysis, the impacts of phase unbalance on DER flexibility aggregation remain largely unexplored. To address these gaps in knowledge, this work proposes a framework for modelling flexibility services in low voltage (LV) distribution networks which enables explicitly imposing voltage unbalance and phase coordination constraints. The simulations, performed for an illustrative 5-bus system and a real 221-bus LV network in the UK, demonstrate that a significant share (over 30%) of total aggregated DER flexibility potential may be unavailable due to voltage unbalances and lack of coordination between DER connected to different phases.
@@ I Introduction
Modern distribution systems are developing with the rapid integration of distributed energy resources (DER), especially in low voltage (LV) networks [1, 2, 3]. Controllable DER, such as battery energy storage systems, prosumers, electric vehicle aggregators, etc., have the technical ability to adjust their power exchange with the grid and, therefore, can provide flexibility services, e.g., by balancing electricity supply and demand [4, 5]. These services can be used to improve the planning and operation of distribution systems, as well as be aggregated and traded between distribution and transmission system operators (DSO and TSO) [6, 7, 8].
@@
Flexibility aggregation models are considered key for providing both local flexibility services (e.g., within LV feeders) and system-level services (e.g., at the primary substations or TSO-DSO interfaces) [9, 10, 11, 12, 13, 3, 14, 5, 15, 16, 17, 18, 19, 20]. The idea behind such models is to estimate the limits of flexible active and reactive power exchanges at certain locations of a distribution network. By estimating these limits, DSO can clear flexibility markets while managing network constraints. The sets of all feasible power exchanges (operating points) in the P-Q space are known as aggregated P-Q flexibility areas or nodal operating envelopes [5, 14].
Despite the advances in DER flexibility aggregation models, there are several critical assumptions made in the literature regarding DER operation and coordination in LV networks. Specifically, this work focuses on the following two problems:
@@
  • Problem #1: Existing studies on DER flexibility aggregation and P-Q flexibility areas calculation assume that all flexibility providers (referred to as flexible units in this work) can be perfectly coordinated [9, 10, 11, 12, 5, 15, 16, 17, 18, 19, 20]. This assumption stems from the use of centralised aggregation methods, where units at different locations can jointly adjust their power generation and consumption in order to manage voltage constraints and keep the network’s operation feasible. However, in real distribution systems, flexible units connected to different phases across the network may not be perfectly coordinated due to low observability, data exchange issues, single-phase connection, behind-the-meter devices, and other constraints, as discussed in [14, 13]. Therefore, there is a problem of incorrect estimation of flexible capacity in distribution networks due to insufficient DER coordination, which can be exacerbated by the next problem.
  • Problem #2: There is a lack of understanding of potential phase unbalance impacts on flexibility services provision. This is partly due to the fact that many existing studies on TSO-DSO flexibility markets rely on single-phase power flow models to simulate flexibility in distribution systems, thus assuming that they are perfectly balanced [9, 10, 11, 12, 5, 14, 15, 16, 17, 18, 19, 20]. Other works that do consider the unbalanced nature of distribution networks, e.g., [13, 3], use three-phase optimal power flow (OPF) models for analysing flexibility in LV distribution networks. Such models are capable of managing network constraints across multiple phases and ensuring the feasibility of DER operation in unbalanced networks. However, they do not specifically assess the impacts of phase unbalance on flexibility services, e.g., the impact of voltage unbalance on aggregated P-Q flexibility limits. Considering the coordination problem between flexible units, phase unbalance can become a major bottleneck in aggregating DER flexibility.
@@
These problems are present in real distribution systems as LV distribution networks are inherently unbalanced and the increase of single-phase uncoordinated distributed generation will further aggravate existing voltage unbalances [21, 22]. For example, suboptimal selection of the DER connection phase can lead to voltage unbalance and limit DER capacity [23]. The need for simulating DER flexibility using accurate three-phase power flow models is acknowledged by the literature [24, 2, 25, 3, 26]. However, existing studies that exploit three-phase power flow models focus on assessing the benefits of DER flexibility services or characterising operating envelopes (limits) for prosumers. Despite recent work [27], there has been no attempt to analyse the aggregated P-Q flexibility for unbalanced LV distribution networks. Note that although P-Q flexibility areas were calculated using a three-phase OPF model in [27], the impact of voltage unbalance constraints was not explicitly quantified. Moreover, no studies have investigated the loss of coordination between DER connected to different phases and its consequences for voltage unbalance and the feasibility of flexibility services provision.
@@
In this regard, this paper investigates the provision of flexibility services in unbalanced LV distribution networks and proposes a framework for quantifying the impacts of phase unbalance and imperfect coordination between distributed flexibility providers. At the core of the framework lies an accurate nonlinear three-phase AC OPF model designed for simulating the operation of flexible units connected to different phases. To characterise feasible flexibility services available in distribution networks, the concept of aggregated P-Q flexibility areas is applied. Then, voltage unbalance constraints are explicitly included to limit the deviations in phase voltages caused by the provision of flexibility services. Finally, phase coordination constraints are imposed to simulate situations where flexible units connected to different phases cannot perform coordinated power adjustment to maximise flexibility services and keep voltage unbalance at acceptable levels. The combination of these models and concepts allows to directly translate voltage unbalance and phase coordination constraints as reductions of the aggregated P-Q flexibility.
@@
To the best of the authors’ knowledge, this work is the first detailed analysis of phase unbalance and DER coordination impacts on aggregated P-Q flexibility in LV distribution networks. Specifically, the paper makes the following contributions:
  • Contribution #1: A novel framework is proposed for quantifying the impacts of phase unbalance and DER coordination on aggregated flexibility in LV distribution networks. The framework comprises three-phase AC power flow models, the concept of P-Q flexibility areas, and voltage unbalance constraints. The combination of these models and concepts allows to directly translate voltage unbalance and phase coordination constraints as reductions in the aggregated P-Q flexibility.
  • Contribution #2: New quantification analysis is performed to demonstrate that phase unbalance and lack of coordination between DER connected to different phases can lead to significant losses in potential aggregated DER flexibility. Simulations performed for an illustrative 5-bus system and a real 221-bus LV network in the UK show that over 30% of total aggregated DER flexibility potential may be unavailable.
Moreover, through extensive simulations for case studies with different combinations of voltage unbalance and phase coordination constraints, the following findings have been formulated:
@@
  • Finding #2: Several small single-phase flexible units have the potential to provide more flexible power than a few larger three-phase balanced flexible units. However, perfect coordination of single-phase units is required to achieve this potential. Without coordination, single-phase flexible units can exacerbate existing voltage unbalances.
  • Finding #3: The worst conditions for providing flexibility services include the lack of coordination between flexible units connected to different phases and tight voltage unbalance constraints. In such cases, flexible units cannot effectively manage voltage unbalance across different phases and locations, which results in the infeasibility of services provision.
@@ II Modelling Framework
To accurately simulate LV distribution networks with flexible units connected to different phases, a multi-phase flexibility estimation model (1a)-(1p) is formulated as a nonlinear programming problem (NLP). The formulation is based on the bus injection model widely applied in unbalanced distribution network optimisation studies [29, 30]. Variables of the model are defined for each phase $\phi$ and include complex voltages and complex power injections at each bus $i$, $V_{i,\phi}$, $s_{i,\phi}$, complex power flows for each line $(i,j)$, $s_{ij,\phi}$, active and reactive power of generators $P_{g,i,\phi}$, $Q_{g,i,\phi}$, and active and reactive power of flexible units $P_{f,i,\phi}$, $Q_{f,i,\phi}$. Voltage unbalances for the selected set of buses $\mathcal{N}^{\text{vu}}$ are explicitly modelled using the positive and negative voltage sequences, ${V_{1}}_{i},{V_{2}}_{i}$, and voltage unbalance factors (VUF), $VUF_{i}$. Note that the original mathematical formulation of the three-phase power flow problem is adapted from the package PowerModelsDistribution.jl [30]. It is then extended by including flexible units, voltage unbalance limits, and phase coordination constraints.
The objective function (1a) minimises or maximises the network’s active and reactive power injections at the reference bus, where coefficients $\alpha^{p}_{i}$ and $\alpha^{q}_{i}$ are introduced to control the optimisation directions in the P-Q space. In this formulation, reference bus $i^{\text{ref}}$ is the location for which flexibility from multiple flexible units is aggregated. Thus, the objective function with adjustable coefficients allows to iteratively (point by point) estimate the limits of aggregated flexibility in the P-Q space at the reference bus [14, 31]. Equations (1b)-(1k) represent the nonlinear AC OPF constraints. Specifically, voltage limits are imposed in (1b), complex power flows and bus injections are defined in (1c), (1d), where $Y_{ij}$ is the admittance matrix. Active and reactive power balance constraints for each node are introduced in (1e), (1f), where $d\in\mathcal{D}$ represents loads of the network. Line capacity limits are included in (1g). Active and reactive power of generators and flexible units is limited in (1h), (1i) and (1j), (1k). The formulated AC OPF model (1a)-(1k) enables estimating the limits of flexibility services for the selected phase and bus of a distribution network, subject to available flexible resources and network constraints.
@@
$\min\enskip\alpha^{p}_{i}\Re(s_{i,\phi})+\alpha^{q}_{i}\Im(s_{i,\phi})\quad$ $i=i^{\text{ref}}\quad\vspace{1\jot}$ (1a)
@@
$Q_{f,i,\phi}^{\min}\leq Q_{f,i,\phi}\leq Q_{f,i,\phi}^{\max}$ $\forall f\in\mathcal{F}\vspace{1\jot}$ (1k)
@@
$VUF_{i}\leq\overline{VUF}$ $i\in\mathcal{N}^{\text{vu}}\vspace{1\jot}$ (1o)
@@
$\sum_{i\in\mathcal{N},f\in\mathcal{F}}P_{f,i,\varphi}=0,\enskip\sum_{i\in\mathcal{N},f\in\mathcal{F}}Q_{f,i,\varphi}=0$ $\forall\varphi\neq\phi\vspace{-1\jot}$ (1p)
@@
These inequalities fix P or Q components of the aggregated flexibility (with a negligibly small margin $\epsilon$) to specific values of interval $k$, $p_{i,\phi}^{k}$ or $q_{i,\phi}^{k}$. Then, model (1a)-(1p) can be solved twice to find the maximum and minimum of the non-fixed flexibility component for this interval. By iteratively imposing the $\epsilon$ constraints and solving model (1a)-(1p), the boundary of the P-Q flexibility area can be calculated as a piece-wise linear approximation with $2k$ operating points. This method gives an easy control over the accuracy (granularity) of the aggregated flexibility estimation. With larger $k$, the accuracy of P-Q flexibility area estimation increases, but the computational cost also grows as the higher numbers of AC OPF problems have to be solved.
Figure 1: Flowchart of the proposed multi-phase flexibility quantification framework: flexibility services from DER are estimated using the concept of P-Q flexibility areas; then, the impact of phase unbalance and DER coordination is analysed by introducing the corresponding constraints.
Note that existing studies on TSO-DSO flexibility markets estimate the P-Q flexibility areas under the assumption of balanced power flow. In this work, the concept of flexibility areas is extended to a multi-phase context: flexibility service limits are estimated in the P-Q space individually for each phase. This requires selecting a reference phase $\phi^{\text{ref}}$ for which the limits of P-Q flexibility will be calculated. The multi-phase formulation enables quantifying the impacts of voltage unbalance and phase coordination constraints. A high-level overview of the proposed framework is presented in Fig. 1. The inputs include network data, available flexible resources, load and flexibility unbalances, VUF limit, and phase coordination assumptions. Then, the limits of flexibility services from DER are estimated using the concept of P-Q flexibility areas. Next, voltage unbalance and phase coordination constraints are imposed and the P-Q flexibility areas are estimated again. Finally, the aggregated flexibility is compared for cases with and without these constraints and the impacts of phase unbalance and DER coordination are quantified (as reductions in the aggregated P-Q flexibility areas, e.g., in kVAr2).

### v2

# Exposing Barriers to Flexibility Aggregation in Unbalanced Distribution Networks

@@ Abstract
The increasing integration of distributed energy resources (DER) offers new opportunities for distribution system operators (DSO) to improve network operation through flexibility services. To utilise flexible resources, various DER flexibility aggregation methods have been proposed, such as the concept of aggregated P-Q flexibility areas. Yet, many existing studies assume perfect coordination among DER and rely on single-phase power flow analysis, thus overlooking barriers to flexibility aggregation in real unbalanced systems. To quantify the impact of these barriers, this paper proposes a three-phase optimal power flow (OPF) framework for P-Q flexibility assessment, implemented as an open-source Julia tool 3FlexAnalyser.jl. The framework explicitly accounts for voltage unbalance and imperfect coordination among DER in low voltage (LV) distribution networks. Simulations on an illustrative 5-bus system and a real 221-bus LV network in the UK reveal that over 30% of the theoretical aggregated flexibility potential can be lost due to phase unbalance and lack of coordination across phases. These findings highlight the need for improved flexibility aggregation tools applicable to real unbalanced distribution networks.
@@ I Introduction
Modern distribution systems are evolving rapidly with the increasing integration of distributed energy resources (DER), particularly in low voltage (LV) networks [1, 2, 3]. Controllable DER, such as battery energy storage systems, prosumers, and electric vehicle aggregators, have the technical ability to adjust their power exchange with the grid and thus provide flexibility services, for example, by helping to balance electricity supply and demand [4, 5]. Such services can improve the operation of distribution systems and can also be aggregated and traded between distribution and transmission system operators (DSO and TSO) [6, 7, 8].
@@
To enable flexibility services, flexibility aggregation models are essential both at the local level (within LV feeders) and system level (at primary substations or TSO-DSO interfaces) [9, 10, 11, 12, 13, 3, 14, 5, 15, 16, 17, 18, 19, 20]. These models aim to estimate the limits of flexible active (P) and reactive (Q) power exchanges at specific locations of a distribution network. By estimating the P-Q limits, DSO can clear flexibility markets while managing network constraints. The sets of all feasible power exchanges (operating points) in the P-Q space are known as aggregated P-Q flexibility areas or nodal operating envelopes [5, 14].
Despite recent advances in DER flexibility aggregation models, several critical assumptions regarding DER operation in unbalanced LV networks remain largely unchallenged, making it easy to overlook barriers to flexibility aggregation in real-world distribution networks. Specifically, this paper exposes two barriers that can significantly reduce the usable flexibility from DER:
@@
  • Barrier #1: Lack of DER coordination. Most flexibility aggregation models assume that all flexibility providers (referred to as flexible units in this work) can be perfectly coordinated [9, 10, 11, 12, 5, 15, 16, 17, 18, 19, 20]. Such centralised aggregation methods imply that units at different locations can jointly adjust their power in order to manage voltage constraints and keep the network’s operation feasible. However, in real distribution systems, flexible units connected to different phases across the network may not be perfectly coordinated due to low observability, data exchange issues, single-phase connection, behind-the-meter devices, and other constraints [14, 13]. These practical limitations of DER coordination create a barrier to effective flexibility aggregation.
  • Barrier #2: Impacts of phase unbalance. Many existing studies rely on single-phase power flow models that implicitly assume perfectly balanced conditions in distribution networks [9, 10, 11, 12, 5, 14, 15, 16, 17, 18, 19, 20]. Although a few works on flexibility services do consider unbalanced distribution systems and apply three-phase optimal power flow (OPF) models, for example [2, 13, 3], they do not quantify the ensuing voltage unbalances and the volume of flexibility lost due to voltage unbalance limits.
In summary, the impacts of voltage unbalance on aggregated P-Q flexibility areas remain largely unexplored, especially when combined with limited coordination among DER. Together, these factors can pose major barriers to flexibility aggregation, making a significant share of theoretical flexibility practically unusable.
@@
These barriers are present in real distribution systems as LV networks are inherently unbalanced and the increase of single-phase uncoordinated distributed generation will further aggravate existing voltage unbalances [21, 22]. For example, suboptimal selection of DER connection phases can lead to voltage unbalance and limit DER capacity [23, 24]. The need for simulating DER flexibility using accurate three-phase power flow models is acknowledged by the literature [25, 2, 26, 3, 27]. However, aside from the recent work [28], there have been no attempts to analyse the aggregated P-Q flexibility for unbalanced LV distribution networks. Note that although P-Q flexibility areas were calculated using a three-phase OPF model in [28], the impact of voltage unbalance constraints was not explicitly quantified. Moreover, no studies have analysed the loss of coordination between DER connected to different phases and its impacts on voltage unbalance.
@@
In this regard, this paper investigates the provision of flexibility services in unbalanced LV distribution networks and proposes a framework for quantifying the impacts of phase unbalance and imperfect coordination between distributed flexibility providers. At the core of the framework lies an accurate nonlinear three-phase AC OPF model designed for simulating the operation of flexible units connected to different phases. To characterise feasible flexibility services available in distribution networks, the concept of aggregated P-Q flexibility areas is applied. Then, voltage unbalance constraints are explicitly included to limit the deviations in phase voltages caused by flexibility services. Finally, phase coordination constraints are imposed to simulate situations where flexible units connected to different phases are unable to adjust their power outputs in a coordinated manner. In mathematical optimisation terms, these constraints fix the power injections of DER in other phases, effectively removing them as control variables in the optimisation problem. Such constraints simulate the practical challenge of coordinating single-phase DERs connected across unbalanced phases without full observability or control. By combining these modelling elements, the framework directly translates voltage unbalance and phase coordination constraints into reductions in the aggregated P-Q flexibility.
@@
To the best of the authors’ knowledge, this work presents the first detailed analysis of the impacts of phase unbalance and DER coordination on aggregated P-Q flexibility in LV distribution networks. Specifically, the paper makes the following contributions:
  • Contribution #1: A novel framework is proposed for quantifying the impacts of phase unbalance and DER coordination on aggregated flexibility in LV distribution networks. By combining three-phase AC power flow models and the concept of P-Q flexibility areas, the framework enables direct translation of voltage unbalance and phase coordination constraints into reductions in the aggregated P-Q flexibility.
  • Contribution #2: New quantification analysis is performed to demonstrate that phase unbalance and lack of coordination between DER connected to different phases can lead to significant losses in aggregated DER flexibility. Simulations performed for an illustrative 5-bus system and a real 221-bus LV network in the UK show that over 30% of the theoretical aggregated flexibility potential can be practically unusable.
Moreover, extensive simulations for case studies with different combinations of voltage unbalance and phase coordination constraints have yielded the following key findings:
@@
  • Finding #2: Several small single-phase flexible units have the potential to provide more flexible power than a few larger three-phase balanced flexible units. However, perfect coordination of single-phase units is required to achieve this potential.
  • Finding #3: Without coordination, single-phase flexible units can exacerbate existing voltage unbalances. A stochastic analysis shows that introducing uncoordinated flexible resources frequently leads to voltage unbalance violations, occurring in over 15% of simulated scenarios.
  • Finding #4: The worst conditions for providing flexibility services include the lack of coordination between flexible units connected to different phases and tight voltage unbalance constraints. In such cases, flexible units cannot effectively manage voltage unbalance across different phases and locations, which results in the infeasibility of services provision.
@@ II Modelling Framework
To accurately simulate LV distribution networks with flexible units connected to different phases, a multi-phase flexibility estimation model (1a)-(1p) is formulated as a nonlinear programming problem (NLP). The formulation is based on the bus injection model widely applied in unbalanced distribution network optimisation studies [30, 31]. Variables of the model are defined for each phase $\phi$ and include complex voltages and complex power injections at each bus $i$, $V_{i,\phi}$, $s_{i,\phi}$, complex power flows for each line $(i,j)$, $s_{ij,\phi}$, active and reactive power of generators $P_{g,i,\phi}$, $Q_{g,i,\phi}$, and active and reactive power of flexible units $P_{f,i,\phi}$, $Q_{f,i,\phi}$. Note that the outputs of flexible units for each phase are the primary control variables of the optimisation problem used to construct the aggregated P-Q flexibility of the network. Voltage unbalances for the selected set of buses $\mathcal{N}^{\text{vu}}$ are explicitly modelled using the positive and negative voltage sequences, ${V_{1}}_{i},{V_{2}}_{i}$, and voltage unbalance factors (VUF), $VUF_{i}$. The original mathematical formulation of the three-phase power flow problem is adapted from the PowerModelsDistribution.jl package [31]. It is then extended by including flexible units, voltage unbalance limits, and phase coordination constraints.
The objective function (1a) minimises or maximises the network’s active and reactive power injections at the reference bus, where coefficients $\alpha^{p}_{i}$ and $\alpha^{q}_{i}$ are introduced to control the optimisation directions in the P-Q space. In this formulation, reference bus $i^{\text{ref}}$ is the location for which flexibility from multiple flexible units is aggregated. Thus, the objective function with adjustable coefficients $\alpha^{p}_{i}$, $\alpha^{q}_{i}$ allows to iteratively (point by point) estimate the limits of aggregated flexibility in the P-Q space at the reference bus [14, 32]. The role of the objective function (1a) in algorithms for constructing P-Q flexibility areas is further discussed later in this section, with a step-by-step illustration provided in Fig. 1. Equations (1b)-(1k) represent the nonlinear AC OPF constraints. Specifically, voltage limits are imposed in (1b), complex power flows and bus injections are defined in (1c), (1d), where $Y_{ij}$ is the admittance matrix. Active and reactive power balance constraints for each node are introduced in (1e), (1f), where $d\in\mathcal{D}$ represents loads of the network. Line capacity limits are included in (1g). Finally, active and reactive power of generators and flexible units is limited in (1h)-(1k).
@@
$\min\enskip\alpha^{p}_{i}\Re(s_{i,\phi})+\alpha^{q}_{i}\Im(s_{i,\phi})\quad$ $i=i^{\text{ref}}\quad\vskip 3.0pt$ (1a)
@@
$Q_{f,i,\phi}^{\min}\leq Q_{f,i,\phi}\leq Q_{f,i,\phi}^{\max}$ $\forall f\in\mathcal{F}\vskip 3.0pt$ (1k)
@@
$VUF_{i}\leq\overline{VUF}$ $i\in\mathcal{N}^{\text{vu}}\vskip 3.0pt$ (1o)
@@
$\sum_{i\in\mathcal{N},f\in\mathcal{F}}P_{f,i,\varphi}=0,\enskip\sum_{i\in\mathcal{N},f\in\mathcal{F}}Q_{f,i,\varphi}=0$ $\forall\varphi\neq\phi\vskip-3.0pt$ (1p)
@@
To investigate the impact of coordination between flexible resources connected to different phases, phase coordination constraints (1p) are introduced. These constraints fix the power injections of DER connected to other phases, effectively removing them as control variables in the optimisation problem. Specifically, in (1p), flexible units not connected to the selected phase $\phi$, that is, $\forall\varphi\neq\phi$, are not allowed to provide flexibility services and support the operation of units connected to phase $\phi$. The inclusion of these constraints models the lack of coordination between flexible units connected to different phases, which can exacerbate voltage unbalance in the network.
@@
Figure 1: Step-by-step illustration of the interval-based OPF algorithm used to construct the aggregated P-Q flexibility area for a single phase: (a) the initial operating point with no flexibility is calculated; (b) the bounds of the area are obtained by calculating Q minimum and maximum, points 1 and 2, and P minimum and maximum, points 3 and 4; (c) these limits are discretised into intervals – 10 Q intervals in this example; (d) the first Q interval is considered by constraining the Q component of the aggregated flexibility, the corresponding P limits (boundary points 5 and 6) are calculated; (e) after evaluating all intervals, the final area is estimated as the hull of all boundary points. The algorithm is executed separately for each phase.
These inequalities fix P or Q components of the aggregated flexibility (with a negligibly small margin $\epsilon$) to specific values of interval $k$, $p_{i,\phi}^{k}$ or $q_{i,\phi}^{k}$. Then, model (1a)-(1p) can be solved twice to find the maximum and minimum of the non-fixed flexibility component for this interval. By iteratively imposing the $\epsilon$ constraints and solving model (1a)-(1p), the boundary of the P-Q flexibility area can be calculated as a piece-wise linear approximation with $2k$ operating points. This method gives an easy control over the accuracy (granularity) of the aggregated flexibility estimation. With larger $k$, the accuracy of P-Q flexibility area estimation increases, but the computational cost also grows as the higher numbers of AC OPF problems have to be solved. Fig. 1 provides a visual overview of the interval-based iterative OPF algorithm for constructing the boundary of P-Q flexibility area.
In this work, the concept of flexibility areas is extended to a multi-phase context: flexibility service limits are estimated in the P-Q space individually for each phase. This requires selecting a reference phase $\phi^{\text{ref}}$ for which the limits of P-Q flexibility will be calculated. The multi-phase formulation enables quantifying the impacts of voltage unbalance and phase coordination constraints. A high-level overview of the proposed framework is presented in Fig. 2. The inputs include network data, available flexible resources, load and flexibility unbalances, VUF limit, and phase coordination assumptions. Then, the limits of flexibility services from DER are estimated using the concept of P-Q flexibility areas. Next, voltage unbalance and phase coordination constraints are imposed and the P-Q flexibility areas are estimated again. Finally, the aggregated flexibility is compared for cases with and without these constraints and the impacts of phase unbalance and DER coordination are quantified (as reductions in the aggregated P-Q flexibility areas, e.g., in kVAr2).
Figure 2: Flowchart of the proposed framework. Flexibility services from DER are estimated for each phase using the concept of P-Q flexibility areas. Then, the impact of phase unbalance and DER coordination is quantified by introducing corresponding constraints, resulting in reduced flexibility areas.